# Bug Fixing with Broader Context: Enhancing LLM-Based Program Repair via Layered Knowledge Injection

### v1

@@ Abstract
Index Terms:
automated program repair, large language models, knowledge injection, in-context learning
@@ I Introduction
  • Replication package [^1] To be released with our bug-type annotations, prompt designs, and evaluation pipeline to support reproducibility and future work on context-aware APR with LLMs.
@@ III-C Research Questions and Information Types in Each Layer
Co-occurring Files. Prior APR research has shown that leveraging historical context can improve repair outcomes [4, 28]. To identify files related to the one containing the buggy function, we analyze past commits and extract those that have most frequently been committed alongside the buggy file [29]. We refer to these as Co-occurring Files. These files suggest a strong logical or functional connection, indicating that they often work together or depend on one another [30].
Structural Dependencies. Understanding the context of the buggy function is critical for generating effective patches [31, 1, 32]. To capture structural dependencies, we examine both directions: the functions and classes invoked by the buggy function and those that invoke the buggy function elsewhere in the codebase. For each buggy function, we use Python’s AST and Jedi [33] library to parse its body and extract referenced functions and classes. We then use the file’s import statements to resolve where those declarations originate and extract their definitions from the relevant files. In the other direction, we identify functions and classes that invoke the buggy function. Since usages can be widespread and introduce noise, we limit this search to the set of previously identified Co-occurring Files. We then apply the same dependency resolution process to extract only the most relevant calling contexts.
@@
III-D Evaluation Metrics
@@
[…]
@@
#$\#\text{fixed}\ (\text{LLM}(J))\triangleq\left|\{b\mid j\in J,C_{j}>0\}\right|$ (2)
@@ IV-A RQ1: To what extent can LLMs resolve different types of bugs using only immediate bug knowledge?
Repository Knowledge Llama 3.3 0.50 0.63 0.68 74% (235/314)
Layer GPT-4o-mini 0.40 0.54 0.61 70% (221/314)
@@ IV-B RQ2: How does injecting repository-level knowledge enhance LLM-based repair of different bug types?
Discussion. These results highlight the role of repository-level knowledge, such as related files, structural dependencies, recent commit history, in enhancing LLM-based bug repair. When local bug knowledge is insufficient, this additional context allows the model to better understand how the buggy function fits within the broader codebase. The fact that both models improve across bug types suggests that repository knowledge is generically beneficial. Moreover, Llama 3.3 continues to show stronger consistency in this layer, with higher Pass@k scores across all bug types. This suggests that it may be better equipped to take advantage of deeper contextual code information when available compared to GPT-4o-mini.
To better understand how repository-level information contributes to bug repair, we manually analyzed the additional bugs solved in this layer (28 by Llama and 24 by GPT-4o-mini). We found that structural dependencies, related files, and commit history provide crucial missing context that helps LLMs infer the intended behavior of the code and produce accurate fixes. For example (Figure 3), a Program Anomaly bug from pandas [49] required changing “index.values” to “index._values” in one of the lines. In the Bug Knowledge Layer, Llama was unaware of “_values” as a valid attribute due to its absence in the local context. However, the structural dependencies and the latest commit affecting the function added additional relevant information, and Llama was able to understand the correct usage and generate the right patch.
@@
To better understand the impact of the Project Knowledge Layer, we manually analyzed the newly fixed bugs in this layer, 15 for Llama 3.3 and 7 for GPT-4o-mini, and examined the generated patches. For example, in a Program Anomaly in the youtube-dl project [55] (Figure 4), the root cause was a regex pattern failing to handle edge cases like integers beginning with zero. Our documentation queries returned a precise constraint: “The function must recognize and convert integers in both decimal and hexadecimal (prefixed with 0x or 0X) and octal (prefixed with 0)”. Combined with the buggy function and relevant API context, this likely enabled the model to generate correct patches.
@@ V Threats to Validity
Construct Validity. A threat to construct validity might arise from the manual bug type annotations. To mitigate this threat, we used an iterative coding process including multiple rounds of discussion and conflict resolution. Each annotator has 3+ years of experience in programming and qualitative research. We also observed high Cohen’s Kappa inter-rater agreement of 0.8, indicating strong agreement [19].
@@ VI Related Work
Prompt Engineering and In-context Learning for APR. Recent research has explored a wide range of techniques to enhance the effectiveness of LLMs in fixing software bugs, including prompt engineering, fine-tuning, and context injection. Xia et al. [61] first showed that prompting LLMs with buggy functions, without any additional context or customization, outperforms traditional APR tools. Building on this, MMAPR[62] and RING [63] applied prompt learning for syntactic and multi-language repair, while InferFix [64] used few-shot prompting to fix static analysis issues. These early systems focused on local bug context and overlooked broader project-level information. The limitations of LLMs became more apparent with the release of the SWE-Bench dataset [65], which includes real-world bugs from large projects. Initial attempts showed low success rates (around 2%), even when LLMs were provided with the localized buggy functions and asked to generate repairs [65]. Ehsani et al. [27] found that different prompting strategies (e.g., chain-of-thought, tree-of-thought) do not yield substantial improvements in bug resolution tasks, suggesting that merely tweaking prompt structure is not enough. To address this, researchers explored integrating richer context. Fan et al. [66] showed that fault localization signals improve fix accuracy. ChatRepair [23] introduced interactive prompting using failing test names and assertions. Other studies found that including local bug context [1], relevant identifiers [67, 2], and structured cues from stack traces [24, 68] improves repair quality. Contextual cues from bug reports and issue descriptions have also been shown to enhance performance [26], as has incorporating historical signals such as previous blame commits [10]. Building on these ideas, Parasaram et al. [9] proposed MANIPLE, a tool that selects and adds bug-related facts into LLM prompts. Their results showed significant gains over naive prompting, especially when space in the context window was limited.
@@ VII Conclusion and Future Work
We introduced a layered knowledge injection framework for LLM-based APR, addressing key limitations in previous work that focused narrowly on isolated local bug context while ignoring the broader structure of real-world software systems. Our approach iteratively injects bug, repository, and project-level contextual information into prompts, enabling more effective and context-aware repair across a diverse set of bug types. Our results confirm the hypothesis posed in our motivating example: the missing piece in LLM-based APR is a deeper understanding of repository and project context. We observed consistent gains in both #fixed and Pass@k scores across layers for both LLama 3.3 and GPT-4o-mini, with LLama achieving a 79% fix rate, a significant improvement of 23% over prior work [9]. All bug types showed improvement from Bug Knowledge Layer to Repository Knowledge Layer, while only a subset (Program Anomaly, Network, and GUI) further benefited from Project Knowledge Layer.

### v3

# Hierarchical Knowledge Injection for Improving LLM-based Program Repair

@@ I Introduction
  • Evaluation against two baselines: (1) a prior fine-grained fact selection approach by Parasaram et al. [9], and (2) an “all-at-once” baseline where all contextual information is injected simultaneously.
@@
  • Replication package [^1] https://github.com/SOAR-Lab/llm-apr-knowledge-injection with our bug-type annotations, prompt designs, and evaluation pipeline to support reproducibility and future work on context-aware APR with LLMs.
@@ III-C Research Questions and Information Types in Each Layer
Co-occurring Files. Prior APR research has shown that leveraging historical context can improve repair outcomes [4, 28]. To identify files related to the one containing the buggy function, we analyze past commits and extract those that have most frequently been committed alongside the buggy file [29]. We refer to these as Co-occurring Files. These files suggest a strong logical or functional connection, indicating that they often work together or depend on one another [30]. Co-occurring files are not directly used in prompts. We use them as a heuristic to identify where the buggy function may be invoked elsewhere in the codebase, aiding in the discovery of structural dependencies discussed next.
Structural Dependencies. Understanding the context of the buggy function is critical for generating effective patches [31, 1, 32]. To capture structural dependencies, we examine both directions: the functions and classes invoked by the buggy function and those that invoke the buggy function elsewhere in the codebase. For each buggy function, we use Python’s AST and Jedi [33] library to parse its body and extract referenced functions and classes. We then use the file’s import statements to resolve where those declarations originate and extract their definitions from the relevant files. In the other direction, we identify functions and classes that invoke the buggy function. Since usages can be widespread and introduce noise especially for utility functions, we limit this search to the set of previously identified Co-occurring Files. We then apply the same dependency resolution process to extract only the most relevant calling contexts.
@@
III-D Baselines
We compare our approach against two baselines: (1) Parasaram et al. [9]’s approach, which leverages various bug-related contextual information to produce patches; and (2) an approach where information from all three layers (bug, repository, and project knowledge) is injected simultaneously for every bug. The second baseline allows us to assess whether progressive, hierarchical injection is more effective than providing all context at once.
III-E Evaluation Metrics
@@
$\text{Pass@k}=\mathbb{E}_{Q}\left[1-\frac{\binom{n-C}{k}}{\binom{n}{k}}\right$] (1)
@@
$\text{fixed}\ (\text{LLM}(J))\triangleq\left|\{b\mid j\in J,C_{j}>0\}\right|$ (2)
@@ IV-A RQ1: To what extent can LLMs resolve different types of bugs using only immediate bug knowledge?
Repository Llama 3.3 0.50 0.63 0.68 74% (235/314)
Knowledge Layer GPT-4o-mini 0.40 0.54 0.61 70% (221/314)
@@ IV-B RQ2: How does injecting repository-level knowledge enhance LLM-based repair of different bug types?
Discussion. These results highlight the role of repository-level knowledge, such as related files, structural dependencies, and recent commit history, in enhancing LLM-based bug repair. When local bug knowledge is insufficient, this additional context allows the model to better understand how the buggy function fits within the broader codebase. The fact that both models improve across bug types suggests that repository knowledge is generically beneficial. Moreover, Llama 3.3 continues to show stronger consistency in this layer, with higher Pass@k scores across all bug types. This suggests that it may be better equipped to take advantage of deeper contextual code information when available compared to GPT-4o-mini.
To better understand how repository-level information contributes to bug repair, we manually analyzed the additional bugs solved in this layer (28 by Llama and 24 by GPT-4o-mini). For every fixed bug, we examined the structure and logic of the buggy function, the additional context retrieved at the current layer, the generated patch and corresponding test outcomes, and then assessed whether the added context plausibly enabled the fix. Manual analysis was done by the first author, and findings were refined through collaborative thematic analysis with the other authors during group discussions, from which we derive the reported patterns and insights. We found that structural dependencies, related files, and commit history provide crucial missing context that helps LLMs infer the intended behavior of the code and produce accurate fixes. For example (Figure 3), a Program Anomaly bug from pandas [49] required changing “index.values” to “index._values” in one of the lines. In the Bug Knowledge Layer, Llama was unaware of “_values” as a valid attribute due to its absence in the local context. However, the structural dependencies and the latest commit affecting the function added additional relevant information, and Llama was able to understand the correct usage and generate the right patch.
@@
To better understand the impact of the Project Knowledge Layer, we manually analyzed the newly fixed bugs in this layer, 15 for Llama 3.3 and 7 for GPT-4o-mini, and examined the generated patches. Our analysis followed the same methodology outlined in RQ2, examining how the additional project-level context contributed to successful repair. This was done by the first author, and the findings were refined through collaborative discussions with all authors. For example, in a Program Anomaly in the youtube-dl project [55] (Figure 4), the root cause was a regex pattern failing to handle edge cases like integers beginning with zero. Our documentation queries returned a precise constraint: “The function must recognize and convert integers in both decimal and hexadecimal (prefixed with 0x or 0X) and octal (prefixed with 0)”. Combined with the buggy function and relevant API context, this likely enabled the model to generate correct patches.
@@
In addition to our hierarchical injection strategy, we compare our approach with an additional baseline where contextual knowledge from all three layers is injected simultaneously, i.e., at the same time, for each bug. As shown in the last two rows of Table II), both LLMs perform worse for all Pass@k metrics when provided with all the layers’ information at once. In terms of the overall percentage of bugs fixed, Llama 3.3 has a fix rate of 65%, identical to its performance using only the first layer, but inferior to the hierarchical configurations that use the repository knowledge and then the project knowledge. GPT-4o-mini performs substantially worse: it fixes 14% fewer bugs compared to its bug-level configuration. These results suggest that providing all information upfront may prevent the model from distinguishing the most relevant context information for a bug, leading to degraded performance. This further highlights the advantage of our hierarchical approach, which allows the model to solve simpler bugs early using minimal context and reserves broader information for more context-hungry bugs that truly require it.
@@ V Threats to Validity
Construct Validity. A threat to construct validity might arise from the manual bug type annotations. To mitigate this threat, we used an iterative coding process including multiple rounds of discussion and conflict resolution. Each annotator has 3+ years of experience in programming and qualitative research. We also observed a high Cohen’s Kappa inter-rater agreement of 0.8, indicating strong agreement [19].
@@ VI Related Work
Prompt Engineering and In-context Learning for APR. Recent research has explored a wide range of techniques to enhance the effectiveness of LLMs in fixing software bugs, including prompt engineering, fine-tuning, and context injection. Xia et al. [61] first showed that prompting LLMs with buggy functions, without any additional context or customization, outperforms traditional APR tools. Building on this, MMAPR[62] and RING [63] applied prompt learning for syntactic and multi-language repair, while InferFix [64] used few-shot prompting to fix static analysis issues. These early systems focused on local bug context and overlooked broader project-level information. The limitations of LLMs became more apparent with the release of the SWE-Bench dataset [65], which includes real-world bugs from large projects. Initial attempts showed low success rates (around 2%), even when LLMs were provided with the localized buggy functions and asked to generate repairs [65]. Ehsani et al. [27, 66] found that different prompting strategies (e.g., chain-of-thought, tree-of-thought) do not yield substantial improvements in bug resolution tasks, suggesting that merely tweaking prompt structure is not enough. To address this, researchers explored integrating richer context. Fan et al. [67] showed that fault localization signals improve fix accuracy. ChatRepair [23] introduced interactive prompting using failing test names and assertions. Other studies found that including local bug context [1], relevant identifiers [68, 2], and structured cues from stack traces [24, 69] improves repair quality. Contextual cues from bug reports and issue descriptions have also been shown to enhance performance [26], as has incorporating historical signals such as previous blame commits [10]. Building on these ideas, Parasaram et al. [9] proposed MANIPLE, a tool that selects and adds bug-related facts into LLM prompts. Their results showed significant gains over naive prompting, especially when space in the context window was limited.
@@ VII Conclusion and Future Work
We introduced a layered knowledge injection framework for LLM-based APR, addressing key limitations in previous work that focused narrowly on isolated local bug context while ignoring the broader structure of real-world software systems. Our approach iteratively injects bug, repository, and project-level contextual information into prompts, enabling more effective and context-aware repair across a diverse set of bug types. Our results confirm the hypothesis posed in our motivating example: one of the missing pieces in LLM-based APR is a deeper understanding of repository and project context. We observed consistent gains in both #fixed and Pass@k scores across layers for both Llama 3.3 and GPT-4o-mini, with Llama achieving a 79% fix rate, a significant improvement of 23% over prior work [9]. All bug types showed improvement from Bug Knowledge Layer to Repository Knowledge Layer, while only a subset (Program Anomaly, Network, and GUI) further benefited from Project Knowledge Layer. Injecting all contextual information at once underperformed our layered strategy for all bug types, reinforcing the importance of incremental, context-aware knowledge injection for maximizing repair performance.